# TriviaHG: A Dataset for Automatic Hint Generation from Factoid Questions

### v1

@@ 5.3. Automatic Evaluation Method
In this section, we assess the automatic evaluation method by comparing the human-provided scores with the outputs generated by our method for the convergence and familiarity quality attributes.
Figure 9. The Pearson correlation coefficient between human assessments and generated values for the convergence quality. L7, L13, L70 refers to LLaMA 7b, 13b, 70b, respectively.
@@ 5.3.1. Convergence Quality Attribute
As elaborated in Section 4.1, we leverage LLMs as the kernel of the automatic evaluation method to assess the convergence quality attribute of a hint. To this end, we employ LLaMA 7b, LLaMA 13b, LLaMA 70b, GPT 3.5 (Brown et al., 2020b) models, as well as the versions of LLaMA 7b, LLaMA 13b, and LLaMA 70b finetuned on the training set of TriviaHG to examine the accuracy of the evaluation method across different kernels. In addition, as discussed in Section 4.1.3, the computed value of the convergence quality attribute is influenced by the number of candidate answers generated. To determine the optimal model as the kernel and the optimal number of candidate answers, we calculate the Pearson correlation coefficient (Han et al., 2022) between human assessments and automatically generated values for the convergence quality attribute across 2,791 hints. Figure 9 indicates the value of the Pearson correlation coefficient for each number of candidate answers for the various kernels. Based on the observations, we can infer that the evaluation method’s alignment with human assessments for the convergence quality attribute is closely linked to the robustness of the LLM used as the kernel. Notably, the highest correlation is observed with GPT 3.5, followed by finetuned LLaMA 70b and LLaMA 70b vanilla. Furthermore, the findings suggest that finetuned LLaMA on the training set yields evaluations more akin to human judgments compared to LLaMA vanilla models. Moreover, the values indicate that generating $11$ candidate answers yields the closest correlation to human results.
@@ 5.3.2. Familiarity Quality Attribute
As detailed in Section 4.2, we employ the normalized number of page views of the Wikipedia page corresponding to entities as the measure of familiarity quality attribute for hints. For each hint, we explore three distinct methods, including the minimum, average, and maximum. In the minimum (maximum) method, we solely consider the normalized value of an entity that has the minimum (maximum) number of views. Conversely, in the average method, we compute the average of normalized values across all the entities mentioned in the hint. To determine the best approach, we compute the Pearson correlation coefficient and Mean squared error (Han et al., 2022) between human assessments and the automatically generated values for the familiarity quality attribute across 2,791 hints. Table 5 indicates the value of the Pearson correlation coefficient and Mean squared error for all the above approaches. The findings suggest that calculating the average of normalized page view values across entities exhibits the highest correlation with human evaluations.
Table 5. The Pearson correlation coefficient and MSE between human assessments and automatically generated values for the familiarity quality attribute across 2,791 hints.
@@ 5.4. Model Performance
The findings suggest that a robust model, characterized by a greater number of training parameters, produces hints of higher convergence quality. This may be attributed to its capacity to store more knowledge within its parameters, resulting in more accurate hints. Furthermore, the results indicate that finetuned LLaMA models outperform LLaMA vanilla models, demonstrating that finetuning on TriviaHG enhances the hint generation capabilities of LLMs. Moreover, comparing the results between Bing Chat AI and GPT 4 underscores the effectiveness of incorporating retrieved documents alongside stored knowledge. This is exemplified by Bing Chat AI, which utilizes GPT 4 as its kernel.
Table 6. The convergence and familiarity values across 100 questions for different models.
Model Convergence Familiarity
@@ 6. Limitations and Use Cases
  • In this study, our focus was solely on factoid questions. In particular, the answers correspond to named entities, each one having its Wikipedia page.
@@
  • In the convergence quality evaluation method, we found 11 to be the optimal number of candidate answers on TriviaHG. However, this number could vary for other datasets. When applying the evaluation method to other datasets, one could ideally try to determine the optimal number of candidate answers for the particular used data.

### v2

@@ 5.3. Automatic Evaluation Method
We assess the automatic evaluation method by comparing the human-provided scores with the generated HICOS and HIFAS.
Figure 9. The Pearson Correlation Coefficient between human assessments and value of HICOS. L7, L13, L70 refers to LLaMA 7b, 13b, 70b, respectively.
@@ 5.3.1. Convergence Quality Attribute
As elaborated in Section 4.1, we leverage LLMs as the kernel of the automatic evaluation method to assess the convergence quality attribute of a hint (HICOS). To this end, we employ LLaMA 7b, LLaMA 13b, LLaMA 70b, GPT 3.5 (Brown et al., 2020b) models, as well as the versions of LLaMA 7b, LLaMA 13b, and LLaMA 70b finetuned on the training set of TriviaHG to examine the accuracy of the evaluation method across different kernels. In addition, as discussed in Section 4.1.3, the value of HICOS is influenced by the number of candidate answers generated. To determine the optimal model as the kernel and the optimal number of candidate answers, we calculate the Pearson Correlation Coefficient (Han et al., 2022) between human assessments and the value of HICOS across 2,791 hints. Figure 9 indicates the value of the Pearson Correlation Coefficient for each number of candidate answers for the various kernels. Based on the observations, we can infer that the evaluation method’s alignment with human assessments for the convergence quality attribute is closely linked to the robustness of the LLM used as the kernel. Notably, the highest correlation is observed with GPT 3.5, followed by finetuned LLaMA 70b and LLaMA 70b vanilla. Furthermore, the findings suggest that finetuned LLaMA on the training set yields evaluations more akin to human judgments compared to LLaMA vanilla models. Moreover, the values indicate that generating $11$ candidate answers yields the closest correlation to human results.
@@ 5.3.2. Familiarity Quality Attribute
As detailed in Section 4.2, we employ the normalized number of page views of the Wikipedia page corresponding to entities as the measure of familiarity quality attribute for hints (HIFAS). For each hint, we explore three distinct methods, including the minimum, average, and maximum. In the minimum (maximum) method, we solely consider the normalized value of an entity that has the minimum (maximum) number of views. Conversely, in the average method, we compute the average of normalized values across all the entities mentioned in the hint. To determine the best approach, we compute the Pearson Correlation Coefficient and Mean Squared Error (MSE) (Han et al., 2022) between human assessments and the value of HIFAS across 2,791 hints. Table 5 indicates the value of the Pearson Correlation Coefficient and Mean Squared Error for all the above approaches. The findings suggest that calculating the average of normalized page view values across entities exhibits the highest correlation with human evaluations.
Table 5. The Pearson Correlation Coefficient and MSE between human assessments and HIFAS across 2,791 hints.
@@ 5.4. Model Performance
The findings suggest that a robust model, characterized by a greater number of training parameters, produces hints of higher convergence quality. This may be attributed to its capacity to store more knowledge within its parameters, resulting in more accurate hints. Furthermore, the results indicate that finetuned LLaMA models outperform LLaMA vanilla models, demonstrating that finetuning on TriviaHG enhances the hint generation capabilities of LLMs. Moreover, comparing the results between Copilot and GPT 4 underscores the effectiveness of incorporating retrieved documents alongside stored knowledge. This is exemplified by Copilot, which utilizes GPT 4 as its kernel.
Table 6. The HICOS and HIFAS values across 100 questions for different models.
Model HICOS HIFAS
@@ 6. Limitations and Use Cases
  • Our focus was solely on factoid questions, where the answers are named entities, each with its own Wikipedia page.
@@
  • In the convergence quality evaluation method (HICOS), we found 11 to be the optimal number of candidate answers on TriviaHG. However, this number could vary for other datasets. When applying the evaluation method to other datasets, one could ideally try to determine the optimal number of candidate answers for the particular used data.